# Hölder regularity of harmonic functions on metric measure spaces

### v2

@@ Abstract
We introduce the Hölder regularity condition for harmonic functions on metric measure spaces and prove that under mild volume regular condition and upper heat kernel estimate, the Hölder regularity condition, the weak Bakry-Émery non-negative curvature condition, the heat kernel Hölder continuity with or without exponential terms and the heat kernel near-diagonal lower bound are equivalent. As applications, firstly, we prove the validity of the so-called generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open by Devyver, Russ, Yang (Int. Math. Res. Not. IMRN (2023), no. 18, 15537–15583). Secondly, we prove that two-sided heat kernel estimates alone imply gradient estimate for the heat kernel on strongly recurrent fractal-like cable systems, which improves the main results of the aforementioned paper. Thirdly, we obtain Hölder (Lipschitz) estimate for heat kernel on general metric measure spaces, which extends the classical Li-Yau gradient estimate for heat kernel on Riemannian manifolds.
Date: August 18, 2024MSC2010: 28A80, 35K08Keywords: Hölder regularity, gradient estimate, harmonic function, heat kernel, Sierpiński carpet.Jin Gao was supported by National Natural Science Foundation of China (12271282). Jin Gao acknowledged Jiaxin Hu and Eryan Hu for reading the preliminary draft. Part of the work was carried out while Meng Yang was participating a Mini-Symposia in the 9th European Congress of Mathematics (9ECM) in Seville (Spain) between 15 and 19 July 2024, Meng Yang was very grateful to Fabrice Baudoin and Li Chen for the invitation to the symposia.
@@ 1 Introduction
Then any weak solution $u\in W^{1,2}_{loc}(D)$ of $Lu=0$ in $D$, or ($L$-)harmonic function $u\in W^{1,2}_{loc}(D)$, is locally Hölder continuous in $D$ or $u\in C^{\alpha}(D)$, where $\alpha\in(0,1)$ is the Hölder exponent. One only needs the measurability of the coefficients $a_{ij}$ to obtain the Hölder regularity of the weak solution $u$. To prove this theorem, very powerful iteration techniques were developed, including de Giorgi iteration [24] and Moser iteration [49], and were applied in many areas later, see [18, 36, 21, 32, 13, 29] and references therein for analysis and heat kernel on metric measure spaces. However, iteration techniques give the existence of some Hölder exponent $\alpha$ which depends only on $d$, $\lambda$, $\Lambda$, but may not give an explicit value of $\alpha$.
The main purpose of this paper is to consider Hölder regularity of harmonic functions on metric measure spaces, where certain explicit Hölder exponent can be obtained. These metric measure spaces will originate from certain strongly recurrent fractals.
Fractals are new examples with very different phenomena from Riemannian manifolds. On a complete non-compact Riemannian manifold, it was independently discovered by Grigor’yan [27] and Saloff-Coste [50, 51] that the following two-sided Gaussian estimate of the heat kernel
@@
$\frac{C_{1}}{V(x,t^{1/\beta})}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,t^{1/\beta})}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right),$
where $\beta$ is a new parameter called the walk dimension, which is always strictly greater than 2 on fractals. Moreover, we say that a fractal is strongly recurrent if its Hausdorff dimension $\alpha$ is strictly less than its walk dimension $\beta$. For example, on the Sierpiński gasket (see Figure 2), we have $\alpha=\log 3/\log 2<\beta=\log 5/\log 2$, see [14, 41], on the Sierpiński carpet (see Figure 2), we have $\alpha=\log 8/\log 3<\beta\approx 2.09697$, see [6, 7, 11, 8, 35, 46].
@@
Another purpose of this paper is to improve some recent results by Devyver, Russ and the second named author [25], where gradient estimate for heat kernel was considered. Gradient estimate for heat kernel plays an important role in the $L^{p}$-boundedness of the Riesz transform for $p>2$, see [3, 22]. To obtain pointwise upper sub-Gaussian estimate for the gradient of the heat kernel on fractal-like cable systems, they introduced the so-called generalized reverse Hölder inequality and verified this condition on the Vicsek and the Sierpiński gasket cable systems, where the proof can be generalized to the cable systems corresponding to a class of p.c.f. self-similar sets. Since the harmonic extension algorithm was intrinsically needed, their proof can not apply on the Sierpiński carpet cable system. In this paper, we introduce the Hölder regularity condition, which is stronger than the generalized reverse Hölder inequality, and, a little bit surprisingly, can be ensured by pointwise two-sided heat kernel estimates. Therefore, on the one hand, we prove the validity of the generalized reverse Hölder inequality on the Sierpiński carpet cable system, which was left open in [25, Page 15545, Line 2]. On the other hand, to obtain gradient estimate for the heat kernel, we do not need to assume gradient type conditions, such as the generalized reverse Hölder inequality, but only pointwise two-sided heat kernel estimates and some mild volume condition are enough, thus we improve main results of [25, Theorem 1.1, Corollary 1.3].
@@
The next purpose of this paper is to generalize the recent result by Baudoin, Chen [15, Theorem 3.3] about Lipschitz estimate for heat kernel on the unbounded Vicsek set. Here we obtain Hölder (Lipschitz) estimate for heat kernel on general unbounded metric measure spaces, which may behave differently at small scale and large scale. We follow a concise analytical approach, thus avoid to going into the probabilistic framework of fractional diffusions by Barlow [5].
@@
To end the Introduction, let us state briefly our main result, more detailed statement will be given in Section 2. We introduce firstly two scaling functions. Let $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\in(0,+\infty)$ satisfy
$22\leq\beta_{i}\leq\alpha_{i}+1,\alpha_{i}<\beta_{i}\text{ for }i=1,$ 2 .
Let $\Phi,\Psi:(0,+\infty)\to(0,+\infty)$ be given by
$\Phi(r)$ $=r^{\alpha_{1}}1_{(0,1)}(r)+r^{\alpha_{2}}1_{[1,+\infty)}(r),$
$\Psi(r)$ $=r^{\beta_{1}}1_{(0,1)}(r)+r^{\beta_{2}}1_{[1,+\infty)}(r).$
@@
In an unbounded metric measure Dirichlet space, we introduce the Hölder regularity condition HR($\Phi,\Psi$) and prove that under mild volume regular condition V($\Phi$) and upper heat kernel estimate UHK($\Psi$), the following conditions are indeed equivalent.
@@
Since we do not need to assume the existence of any gradient operator associated with the metric measure Dirichlet space, in particular, a “carré du champ”, the above equivalence generalizes the corresponding result by Coulhon, Jiang, Koskela, Sikora [22, Theorem 1.2].
@@
In addition, we also extend the classical result on Riemannian manifolds to metric measure Dirichlet spaces, that is, the conjunction of UHK($\Psi$) and $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ implies NLE($\Psi$) on metric measure Dirichlet spaces, which extends that the conjunction of upper and gradient estimates of heat kernel implies lower estimate of heat kernel on Riemannian manifolds, see Cheeger, Yau [19] and Li, Yau [47].
@@ 2 Statement of the main results
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded metric measure Dirichlet (MMD) space, that is, $(X,d)$ is a locally compact separable unbounded metric space, $m$ is a positive Radon measure on $X$ with full support, $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Throughout this paper, we always assume that all metric balls are relatively compact and $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative.
For any $x\in X$, for any $r\in(0,+\infty)$, denote the (metric) ball $B(x,r)=\{y\in X:d(x,y)<r\}$, denote $V(x,r)=m(B(x,r))$. If $B=B(x,r)$, then we denote $\delta B=B(x,\delta r)$ for any $\delta\in(0,+\infty)$. Let $C(X)$ denote the space of all real-valued continuous functions on $X$ and let $C_{c}(X)$ denote the space of all real-valued continuous functions on $X$ with compact support.
Consider the strongly local regular Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$. Let $\Delta$ be the corresponding generator which is a non-negative definite self-adjoint operator. Let $\Gamma$ be the corresponding energy measure. Denote $\mathcal{E}_{1}(\cdot,\cdot)=\mathcal{E}(\cdot,\cdot)+(\cdot,\cdot)$, where (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is the inner product in $L^{2}(X;m)$. We refer to [26] for related results about Dirichlet forms.
We need to introduce some conditions to state our main results.
@@
We say that the heat kernel upper bound UHK($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C_{1}$, $C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$, we have
@@
If a lower bound, similar to UHK($\Psi$), with different constants $C_{i}$ also holds, then we say that HK($\Psi$) holds.
We say that the heat kernel near-diagonal lower bound NLE($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C,\varepsilon\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$ with $d(x,y)<\varepsilon\Psi^{-1}(t)$, we have
@@
If the metric $d$ is furthermore assumed to satisfy the chain condition, then the conjunction of UHK($\Psi$) and NLE($\Psi$) is equivalent to HK($\Psi$), see [32, Section 6].
Let $D$ be an open subset of $X$. Let
@@
Let $f\in L^{1}_{loc}(D)$. We say that $u\in\mathcal{F}$ is a solution of the Poisson equation $\Delta u=f$ in $D$ if
@@
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$ HR($\Phi,\Psi$)
The above condition is some kind of generalization of the so-called generalized reversed Hölder inequality GRH($\Phi,\Psi$) introduced in [25].
We say that the Hölder estimate HHK($\Phi,\Psi$) for the heat kernel holds if there exists $C_{HHK}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x_{1},x_{2},y_{1},y_{2}\in X$, we have
@@
The above condition implies that for any $t\in(0,+\infty)$, for any $x\in X$, $p_{t}(x,\cdot)$ is $\gamma_{1}$-Hölder continuous in the small scale and $\gamma_{2}$-Hölder continuous in the large scale.
We say that the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel holds if there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
@@
The above condition is some kind of generalization of the weak Bakry-Émery curvature condition introduced in [1, Definition 3.1].
@@
UHK($\Psi$). Then the followings are equivalent.
@@
The key novality of the above result is that the pointwise two-sided heat kernel estimates are enough to give some proper explicit Hölder regularity of harmonic functions. In previous literatures (for example [36, Section 5.3], [21, Proposition 4.5], [32, Section 5.3], [13, Section 4.2.1]), using Moser iteration technique or de Giorgi iteration technique and oscillation inequality, one can obtain only some Hölder exponent depending implicitly on other constants.
Our proof will follow the following diagrams.
@@
In the setting of fractional metric spaces with fractional diffusions introduced in [5, Section 3], when the dimension $\alpha$ of the state space is smaller than the walk dimension $\beta$, or $\alpha_{1}=\alpha_{2}=\alpha<\beta_{1}=\beta_{2}=\beta$ in our setting, it was proved in [1, Lemma 3.4, Remark 3.6, Theorem 3.7] that
@@
In the fractal setting, HHK($\Phi,\Psi$) was also proved on the unbounded Sierpiński gasket [14, Theorem 1.5 (c)] and on the unbounded Sierpiński carpet [8, Theorem 1.1 (e)].
@@
The above equivalence can be regarded as some “Hölder” version generalization of the “gradient” version result [22, Theorem 1.2]. It was proved that in a non-compact doubling Dirichlet metric measure space $(X,d,\mu,\mathscr{E})$ endowed with a “carré du champ”. Assume the upper Gaussian bound $(UE)$ for the heat kernel and a local $L^{\infty}$-Poincaré inequality $(P_{\infty,\text{loc}})$. Then the following conditions are equivalent.
  • The quantitative reverse $L^{\infty}$-Hölder inequality ($\text{RH}_{\infty}$) for gradients of harmonic functions: There exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u$ which is harmonic in $2B$, we have
    $\lVert|\nabla u|\rVert_{L^{\infty}(B)}\leq\frac{C}{r}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}\mu.$ ($\text{RH}_{\infty}$)
  • The pointwise Li-Yau gradient estimate ($\text{GLY}_{\infty}$) for heat kernel: There exist $C,c\in(0,+\infty)$ such that
@@
  • The $L^{\infty}$-boundedness ($\text{G}_{\infty}$) of the gradient of the heat semi-group $|\nabla H_{t}|$: There exists $C\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, we have
@@
Our conditions are “Hölder” version of the above “gradient” conditions as in Figure 3.
“Hölder” conditions “gradient” conditions
@@
Figure 3: The correspondence between “Hölder” and “gradient” conditions
In their proof to obtain ($\text{RH}_{\infty}$), they established a reproducing formula for harmonic functions using the finite propagation speed property, which may not be available on general MMD spaces. We leave it as an open question to generalize the reproducing formula technique to prove HR($\Phi,\Psi$) on MMD spaces or the following GRH($\Phi,\Psi$) on cable systems, in particular, to develop the corresponding finite propagation speed theory.

### v3

@@ Abstract
We introduce a Hölder regularity condition for harmonic functions on metric measure spaces and prove that, under a slow volume regular condition and an upper heat kernel estimate, the Hölder regularity condition, the weak Bakry-Émery non-negative curvature condition, Hölder continuity of the heat kernel (with or without exponential terms), and the near-diagonal lower bound for the heat kernel are equivalent. As applications, first, we establish the validity of the so-called generalized reverse Hölder inequality on the Sierpiński carpet cable system, resolving an open problem left by Devyver, Russ, Yang (Int. Math. Res. Not. IMRN (2023), no. 18, 15537–15583). Second, we prove that two-sided heat kernel estimates alone imply gradient estimates for the heat kernel on strongly recurrent fractal-like cable systems, improving the main results of the aforementioned paper. Third, we obtain Hölder (Lipschitz) estimates for the heat kernel on strongly recurrent metric measure spaces, extending the classical Li-Yau gradient estimate for the heat kernel on Riemannian manifolds.
Date: MSC2020: 28A80, 35K08Keywords: Hölder regularity, gradient estimate, harmonic function, heat kernel, Sierpiński carpet.Jin Gao was supported by National Natural Science Foundation of China (Grant. No. 12271282), and Zhejiang Provincial Natural Science Foundation of China (Grant. No. LQN25A010019). Jin Gao acknowledges Jiaxin Hu and Eryan Hu for reading the preliminary draft. Part of the work was carried out while Meng Yang was participating a Mini-Symposia in the 9th European Congress of Mathematics (9ECM) in Seville (Spain) between 15 and 19 July 2024, Meng Yang is very grateful to Fabrice Baudoin and Li Chen for the invitation to the symposia. The authors are grateful to Fabrice Baudoin for discussions on curvature conditions, and warmly thank Naotaka Kajino and Ryosuke Shimizu for pointing out an alternative proof of Proposition 8.12. The authors thank the anonymous referees for their helpful comments and suggestions, which have improved the presentation of this paper.
@@ 1 Introduction
Then any weak solution $u\in W^{1,2}_{\mathrm{loc}}(D)$ of $Lu=0$ in $D$, that is, any ($L$-)harmonic function $u\in W^{1,2}_{\mathrm{loc}}(D)$, is locally Hölder continuous in $D$; equivalently, $u\in C^{\alpha}(D)$ for some $\alpha\in(0,1)$, where $\alpha$ is a Hölder exponent. One only needs the measurability of the coefficients $a_{ij}$ to obtain Hölder regularity of the weak solution $u$. To prove this theorem, very powerful iteration techniques were developed, including the De Giorgi iteration [25] and the Moser iteration [51], and were later applied in many areas; see [19, 38, 22, 34, 14, 30] and the references therein for applications to analysis and heat kernel estimates on metric measure spaces. However, these iteration techniques give the existence of some Hölder exponent $\alpha$, which depends only on $d,\lambda,\Lambda$, but generally do not provide an explicit value of $\alpha$.
The main purpose of this paper is to consider Hölder regularity of harmonic functions on metric measure spaces, where an explicit Hölder exponent can be obtained. These metric measure spaces arise from certain strongly recurrent fractals.
Fractals provide new examples exhibiting phenomena markedly different from those on Riemannian manifolds. On a complete non-compact Riemannian manifold, it was independently discovered by Grigor’yan [28] and Saloff-Coste [54, 55] that the following two-sided Gaussian estimate of the heat kernel
@@
$\frac{C_{1}}{V(x,t^{1/\beta})}\exp\left(-C_{2}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right)\leq p_{t}(x,y)\leq\frac{C_{3}}{V(x,t^{1/\beta})}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta}}\right)^{\frac{\beta}{\beta-1}}\right),$ $\textrm{HK}(\beta)$
where $\beta$ is a new parameter called the walk dimension, which is typically strictly greater than 2 on fractals. Notably, for $\beta=2$, $\textrm{HK}(\beta)$ is indeed HK(2). For example, on the Sierpiński gasket (see Figure 2), we have $\beta=\log 5/\log 2$, see [15, 43], while on the Sierpiński carpet (see Figure 2), we have $\beta\approx 2.09697$, see [7, 8, 12, 9, 37, 48].
@@
Another purpose of this paper is to improve some recent results of Devyver, Russ and the second named author [26], where gradient estimates for the heat kernel were considered. Gradient estimates for the heat kernel play an important role in the $L^{p}$-boundedness of the Riesz transform for $p>2$; see [4, 23]. To obtain pointwise upper sub-Gaussian estimates for the gradient of the heat kernel on fractal-like cable systems, the authors of [26] introduced the so-called generalized reverse Hölder inequality and verified this condition on the Vicsek and Sierpiński gasket cable systems, where the proof can be extended to cable systems corresponding to a class of p.c.f. self-similar sets. Since the harmonic extension algorithm is intrinsically needed, their proof does not apply to the Sierpiński carpet cable system. In this paper, we introduce a Hölder regularity condition, which is stronger than the generalized reverse Hölder inequality and, somewhat surprisingly, can be ensured by pointwise two-sided heat kernel estimates. As consequences, on the one hand, we establish the validity of the generalized reverse Hölder inequality on the Sierpiński carpet cable system, resolving an open problem left in [26, Page 15545, Line 2]. On the other hand, to obtain gradient estimates for the heat kernel, we do not need to assume gradient-type conditions such as the generalized reverse Hölder inequality; instead, pointwise two-sided heat kernel estimates together with a slow volume regular condition suffice. This leads to an improvement of the main results of [26, Theorem 1.1, Corollary 1.3].
@@
The next purpose of this paper is to generalize a recent result of Baudoin, Chen [16, Theorem 3.3] about Lipschitz estimates for the heat kernel on the unbounded Vicsek set. In this work, we obtain Hölder (Lipschitz) estimates for the heat kernel on strongly recurrent unbounded metric measure spaces, which may exhibit different behaviors at small and large scales. We follow a concise analytic approach, thereby avoiding the probabilistic framework of fractional diffusions introduced by Barlow [6].
@@
We conclude the Introduction by briefly stating our main result; a more detailed statement will be given in Section 2.
Let $(X,d,m)$ be a metric measure space and let $(\mathcal{E},\mathcal{F})$ be a strongly local regular Dirichlet form on $L^{2}(X;m)$. A simplified consequence of our results is that, if $X$ is $\alpha$-regular for some $\alpha>0$, that is, for any metric ball $B(x,r)$, we have
$V(x,r)=m(B(x,r))\asymp r^{\alpha},$
and if $(\mathcal{E},\mathcal{F})$ admits a heat kernel $p_{t}(x,y)$ satisfying $\textrm{HK}(\beta)$ with walk dimension $\beta>\alpha$, then all harmonic functions associated with $(\mathcal{E},\mathcal{F})$ are $(\beta-\alpha)$-Hölder continuous. Moreover, the same Hölder regularity holds for the heat kernel. This result is motivated by classical examples on fractals, including the Sierpiński gasket and the Sierpiński carpet; see [15, 9].
In fact, we consider more general and natural assumptions. We assume the volume regular condition
$V(x,r)\asymp\Phi(r),$
where
$\Phi(r)=r^{\alpha_{1}}1_{(0,1)}(r)+r^{\alpha_{2}}1_{[1,+\infty)}(r),$ (1.1)
and $\alpha_{1},\alpha_{2}>0$. $\textrm{HK}(\beta)$ is also generalized by introducing a time-distance scaling function $\Psi$ of the form
$\Psi(r)=r^{\beta_{1}}1_{(0,1)}(r)+r^{\beta_{2}}1_{[1,+\infty)}(r),$ (1.2)
which involves two walk dimensions: $\beta_{1}$ at small scales and $\beta_{2}$ at large scales. We further assume the strongly recurrent condition
$\alpha_{i}<\beta_{i}\text{ for }i=1,2.$ (1.3)
Under these assumptions, the Hölder regularity result is stated with two Hölder exponents: $\beta_{1}-\alpha_{1}$ at small scales and $\beta_{2}-\alpha_{2}$ at large scales.
@@
More precisely, in an unbounded metric measure Dirichlet space, we introduce the Hölder regularity condition HR($\Phi,\Psi$) and prove that, assuming (1.3), the volume regular condition V($\Phi$), and the upper heat kernel estimate UHK($\Psi$), the following conditions are equivalent:
@@
Since we do not assume the existence of any gradient operator associated with the metric measure Dirichlet space, in particular, a “carré du champ”, the above equivalence generalizes the corresponding result of Coulhon, Jiang, Koskela, Sikora [23, Theorem 1.2].
@@
In addition, we extend a classical result from the setting of Riemannian manifolds to metric measure Dirichlet spaces. We prove that the conjunction of UHK($\Psi$) and $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ implies NLE($\Psi$) on metric measure Dirichlet spaces. This generalizes the fact that, on Riemannian manifolds, the conjunction of upper heat kernel bounds and gradient estimates implies a lower heat kernel bound; see Cheeger, Yau [20] and Li, Yau [49].
@@ 2 Statement of the main results
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded metric measure Dirichlet (MMD) space, that is, $(X,d)$ is a locally compact separable unbounded metric space, $m$ is a positive Radon measure on $X$ with full support, and $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Although the results of this paper also apply to bounded spaces, we focus on the unbounded case, where the space may exhibit different behaviors at small and large scales. Throughout this paper, we always assume that all metric balls are relatively compact, and that $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$ is conservative.
For any $x\in X$ and any $r\in(0,+\infty)$, denote the (metric) ball $B(x,r)=\{y\in X:d(x,y)<r\}$, denote $V(x,r)=m(B(x,r))$. If $B=B(x,r)$, then we denote $\delta B=B(x,\delta r)$ for any $\delta\in(0,+\infty)$. Let $C(X)$ denote the space of all real-valued continuous functions on $X$ and let $C_{c}(X)$ denote the space of all real-valued continuous functions on $X$ with compact support. Denote $\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{A}=\frac{1}{m(A)}\int_{A}$ and $u_{A}=\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{A}u\mathrm{d}m$ for any measurable set $A$ with $m(A)\in(0,+\infty)$ and any function $u$ such that the integral $\int_{A}u\mathrm{d}m$ is well-defined.
Consider the strongly local regular Dirichlet form $(\mathcal{E},\mathcal{F})$ on $L^{2}(X;m)$. Let $\Delta$ be the corresponding generator which is a non-negative definite self-adjoint operator. Let $\Gamma$ be the corresponding energy measure. Denote $\mathcal{E}_{1}(\cdot,\cdot)=\mathcal{E}(\cdot,\cdot)+(\cdot,\cdot)$, where $(\cdot,\cdot)$ is the inner product in $L^{2}(X;m)$. We refer to [27] for related results about Dirichlet forms.
We introduce several conditions in order to state our main results.
@@
We say that the upper heat kernel estimate UHK($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C_{1}$, $C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$, we have
@@
If a corresponding lower bound, analogous to UHK($\Psi$) but possibly with different constants $C_{i}$, also holds, then we say that HK($\Psi$) holds.
We say that the near-diagonal lower bound for the heat kernel NLE($\Psi$) holds if there exist a properly exceptional set $\mathcal{N}$ and $C,\varepsilon\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y\in X\backslash\mathcal{N}$ with $d(x,y)<\varepsilon\Psi^{-1}(t)$, we have
@@
If the metric $d$ is furthermore assumed to satisfy the chain condition, then the conjunction of UHK($\Psi$) and NLE($\Psi$) is equivalent to HK($\Psi$), see [34, Section 6]. Moreover, assuming HK($\Psi$), $d$ satisfies the chain condition; see [52, Theorem 2.11]. If, in addition, V($\Phi$) holds, then $2\leq\beta_{i}\leq\alpha_{i}+1$ for $i=1,2$, see [53, THEOREM 2.1] and [32, Corollary 4.9].
Let $D$ be an open subset of $X$ and
@@
Let $f\in L^{1}_{\mathrm{loc}}(D)$. We say that $u\in\mathcal{F}$ is a solution of the Poisson equation $\Delta u=f$ in $D$ if
@@
$\frac{\Phi(d(x,y))}{\Psi(d(x,y))}|u(x)-u(y)|\leq C_{H}\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$ HR($\Phi,\Psi$)
The above condition can be viewed as a generalization of the so-called generalized reversed Hölder inequality GRH($\Phi,\Psi$) introduced in [26].
We say that the Hölder estimate for the heat kernel HHK($\Phi,\Psi$) holds if there exists $C_{HHK}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x_{1},x_{2},y_{1},y_{2}\in X$, we have
@@
The above condition implies that for any $t\in(0,+\infty)$, for any $x\in X$, $p_{t}(x,\cdot)$ is $\gamma_{1}$-Hölder continuous at small scales and $\gamma_{2}$-Hölder continuous at large scales.
We say that the Hölder estimate with exponential terms for the heat kernel $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ holds if there exist $C_{1},C_{2}\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, for any $x,y_{1},y_{2}\in X$, we have
@@
The above condition can be viewed as a generalization of the weak Bakry-Émery non-negative curvature condition introduced in [2, Definition 3.1]. Roughly speaking, letting $t\to 0+$ in wBE($\Phi,\Psi$) suggests that the space is strongly recurrent.
@@
UHK($\Psi$), where $\Phi$ and $\Psi$ are given by (1.1) and (1.2), respectively, with parameters $\alpha_{i},\beta_{i}$, $i=1,2$, satisfying the strongly recurrent condition (1.3). Then the followings are equivalent.
@@
The key novelty of the above result is that pointwise two-sided heat kernel estimates suffice to give explicit Hölder regularity for harmonic functions. In the existing literature (see, for example, [38, Section 5.3], [22, Proposition 4.5], [34, Section 5.3], and [14, Section 4.2.1]), approaches based on the Moser or De Giorgi iteration, together with oscillation inequalities, typically give only Hölder continuity with an exponent that depends implicitly on other constants.
Our proof will follow the implication diagrams below:
@@
In the setting of fractional metric spaces with fractional diffusions introduced in [6, Section 3], when the dimension $\alpha$ of the state space is strictly smaller than the walk dimension $\beta$, that is, when $\alpha_{1}=\alpha_{2}=\alpha<\beta_{1}=\beta_{2}=\beta$ in our notation, it was proved in [2, Lemma 3.4, Remark 3.6, Theorem 3.7] that
@@
In the fractal setting, HHK($\Phi,\Psi$) was also established on the unbounded Sierpiński gasket (see [15, Theorem 1.5 (c)]) and on the unbounded Sierpiński carpet (see [9, Theorem 1.1 (e)]).
@@
The above equivalence can be regarded as a Hölder-type analogue of the gradient-type result in [23, Theorem 1.2]. In that work, it was proved that on a non-compact doubling Dirichlet metric measure space $(X,d,\mu,\mathscr{E})$ endowed with a “carré du champ”, assuming the upper Gaussian bound $(UE)$ for the heat kernel and a local $L^{\infty}$-Poincaré inequality $(P_{\infty,\text{loc}})$, the following conditions are equivalent.
  • ($\text{RH}_{\infty}$) Quantitative reverse $L^{\infty}$-Hölder inequality for gradients of harmonic functions: There exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u$ which is harmonic in $2B$, we have
    $\lVert{|\nabla u|}\rVert_{L^{\infty}(B)}\leq\frac{C}{r}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}\mu.$ ($\text{RH}_{\infty}$)
  • ($\text{GLY}_{\infty}$) Pointwise Li-Yau gradient estimate for the heat kernel: There exist $C,c\in(0,+\infty)$ such that
@@
  • ($\text{G}_{\infty}$) $L^{\infty}$-boundedness of the gradient of the heat semi-group $|\nabla H_{t}|$: There exists $C\in(0,+\infty)$ such that for any $t\in(0,+\infty)$, we have
@@
  • (GBE) A generalized Bakry-Émery condition: There exist $C,c\in(0,+\infty)$ such that for any $f\in W^{1,2}(X)$, for any $t\in(0,+\infty)$ and $\mu$-a.e. $x\in X$, we have
    $\lvert\nabla H_{t}f(x)\rvert^{2}\leq CH_{ct}(\lvert\nabla f\rvert^{2})(x).$ (GBE)
Our conditions can be viewed as Hölder-type analogues of the above gradient conditions, as illustrated in Figure 3.
Hölder-type conditions gradient-type conditions
@@
Figure 3: The correspondence between Hölder-type and gradient-type conditions
In their proof of ($\text{RH}_{\infty}$), they established a reproducing formula for harmonic functions using the finite propagation speed property, which may not be available on general MMD spaces. We leave it as an open question to extend the reproducing formula technique to prove HR($\Phi,\Psi$) on MMD spaces or the following GRH($\Phi,\Psi$) on cable systems, and in particular to develop a corresponding finite propagation speed theory in these settings.